Potassium hydrogen phthalate is a solid, monoprotic acid frequently used in the laboratory as a primary standard. It has the unwieldy formula of KHC8H4O4. This is often written in shorthand notation as KHP.

If 35.6 mL of a barium hydroxide solution are needed to neutralize 2.89 grams of KHP, what is the molarity of the barium hydroxide solution?

_______ M
1
Expert's answer
2019-03-18T08:13:25-0400

The balanced equation for this reaction is:



The number of moles of KHP is:

"n(KHP)=\frac {m(KHP)} {M(KHP)} = \frac {2.89 g} {204 g/mol}=0.01416 mol"

According to the balanced equation, we need twice less moles of barium hydroxide per each mole of KHP. Then, number of moles of the barium hydroxide is equal to 0.5*0.01416 mol = 0.00708 mol.

Now we can calculate the molarity of a barium hydroxide solution:

"C_M(Ba(OH)_2)=\frac {n(Ba(OH)_2)} {V(Ba(OH)_2)} = \frac {0.00708 mol} {0.0356 L}=0.19897M"


Answer: the molarity of the barium hydroxide solution is 0.19897M.